solve the logarithmic equation 7+3 ln(x)=5 and approximate the results to three decimal places
Answer by rothauserc(4718)

7 +3 ln(x) = 5
note that ln is the natural log and base is e
subtract 7 from both sides of =
3 ln(x) = -2
divide both sides of = by 3
ln(x) = -2/3
x = e^(-2/3) = 1 / e^(2/3) = 0.513417119 approx 0.513
x = 0.513